CURRICULUM VITAE

Name  John Rognes
Home address  Røahagan 1 G

N – 0754 Oslo

Norway

Telephone  (47) 22 85 58 45 (office)

(47) 22 85 43 49 (telefax)

(47) 22 50 80 80 (private)

E–mail  rognes@math.uio.no
Birthdate  April 28th 1966 in Oslo
Citizenship  Norwegian

Academic Education

International Baccalaureate diploma, Berg Higher Secondary School, Oslo, 1981–1984. Result: 45 points of 45 possible.

Cand. mag. degree (Bachelor’s degree), University of Oslo, Faculty of Mathematics and Sciences, 1984–1985. Major subject: Mathematics. Minor subjects: Computer Science and Statistics. Average grade: 1.09 (Laudabilis prae ceteris).

Master of Arts degree in Mathematics, Princeton University, October 1987. Major subjects for general examination: Differential Topology and Algebraic Topology.

Ph.D. degree in Mathematics, Princeton University, June 1990. Thesis advisor: Gunnar Carlsson. Thesis title: “The Rank Filtration in Algebraic K-theory.”

Visiting scholar/exchange Ph.D. student at the University of California at Berkeley, August 1989 to May 1990, in connection with the year in Algebraic Topology at the Mathematical Sciences Research Institute (MSRI).

Ph.D. evaluated as comparable to a Norwegian doctor philosophiae, March 1992.

Faculty development course: Teaching and learning in higher education (one year course in pedagogical methods), University of Oslo, 1995.

Engagements, Posts

Computer programmer at Luleå University, Sweden, 1982.

Assistant teacher in Mathematics, University of Oslo, 1983–1986.

Research and computer work during military service at Norwegian Defense Research Establishment, Kjeller, Norway, 1990–1991.

Member of “Graduiertenkolleg für Mathematik,” Universität Bielefeld, June to September 1991.

Adjunkt (Visiting Researcher) in Mathematics, Aarhus University, September 1991 to May 1992.

Fakultetstipendiat (Post Doc.), Mathematics Department, University of Oslo, January 1992 to August 1994.

Visiting researcher, Sonderforschungsbereich 343, “Diskrete Strukturen in der Mathematik,” Universität Bielefeld, June 1993, December 1994 and December 1997.

Post Doctoral Fellow, year in “Topology and Algebraic K-theory,” Mittag–Leffler institute, Stockholm, January to June 1994.

Førsteamanuensis (Associate Professor), Mathematics Department, University of Oslo, August 1994 to April 1998.

Visting Associate Professor, Mathematics Department, Stanford University, January to March 1996.

Visting Associate Professor, Mathematics Department, University of Chicago, April to May 1996.

Professor (Full Professor), Mathematics Department, University of Oslo, from May 1st 1998.

Research Interests

Algebraic K-theory, homotopy theory, geometric topology, gauge theory, number theory.

Teaching Experience

Assistant Teacher, Calculus (MA 100, MA 001 and Mathematics for Economists), University of Oslo, four semesters 1983–1986, and fall 1996.

Assistant Teacher, Linear Algebra (Mat 10) and Differential Geometry (Geo 1), Aarhus University, fall 1991 and spring 1992.

Lecturing, Algebraic Topology (MA 362), University of Oslo, fall 1992, fall 1993 and fall 1999.

Assistant Teacher, Linear Algebra (MA 104), University of Oslo, fall 1992, fall 1995 and fall 1997.

Lecturing, Algebraic Topology II (MA 422), University of Oslo, spring 1993 (The image of J in algebraic K-theory), spring 1995 (Characteristic classes), spring 1997 (Homotopy theory), spring 1998 (Manifold models for A-theory) and spring 1999 (Elliptic curves and chromatic homotopy theory).

Developed course notes for MA 422, “Bildet av J-homomorfien i stabil homotopi og algebraisk K-teori,” TEX–ed lecture notes, ca. 80 pages, spring 1993.

Developed course notes for MA 422, “Manifold models for A-theory,” spring 1998.

Lecturing, General Topology (MA 245), University of Oslo, fall 1994. Lecturing, Introduction to Number Theory (Ma 366), University of Oslo, spring 1995. Lecturing, Seminar on Four-manifolds and Seiberg–Witten invariants, University of Oslo, fall 1995. Lecturing, Matrix Theory and its Applications (Math 103), Stanford University, Winter Quarter 1996. Lecturing, Calculus (MA 100), University of Oslo, fall 1996. Organized Topology seminar, Univerity of Oslo, most semesters 1994–1999. Lecturing, Linear Algebra (MA 104), University of Oslo, fall 1998.

Graduate Students Supervised

Halvard Fausk, Cand. scient. (Master of Arts) student, spring 1994–June 1995. Thesis title: “The chain rule in functor calculus.” Paul Arne Østvær, Cand. scient. (Master of Arts) student, fall 1995–December 1996. Thesis title: “The Bloch–Lictenbaum spectral sequence and calculations in algebraic Ktheory.” Erik Strand, Cand. scient. (Master of Arts) student, fall 1996–December 1998. Thesis title: “Generalized homological algebra.” Eirik Anonsen, Cand. scient. (Master of Arts) student. Proposed thesis topic: “Ring structures on Moore spectra.” Sverre Lunøe–Nilsen, Cand. scient. (Master of Arts) student. Proposed thesis topic: “The involution on the space of smooth tubes.” Hans Jørgen Riddervold, Cand. scient. (Master of Arts) student. Jo Raabe, Cand. scient. (Master of Arts) student. Paul Arne Østvær, Dr. scient. (Ph.D.) student, January 1997–present.

Post Docs Supervised

Stavros Tsalidis, Ph.D. Brown University, 1997–1999. Morten Brun, Ph.D. Aarhus University, 1998–1999. Iver Ottosen, Ph.D. Aarhus University, 1998. Christian Schlichtkrull, Ph.D. Aarhus University, 1998. Christian Ausoni, Ph.D. Lausanne University, 1998–1999.

Invited Conference Talks “K_{4}(Z) is the trivial group,” invited talk at International Conference on Algebra and Topology, Kazimierz, June 1993.

Invited to speak in the Special Session on Homotopy Theory of the Joint Meeting of the American Mathematical Society and the Deutsche Mathematiker–Vereinigung, Heidelberg, October 1993.

“K_{4}(Z) is the trivial group,” main talk at Conference on Algebraic K-theory at Université Paris 7 in July 1994, preceding the 1994 International Congress of Mathematicians in Zürich.

“Algebraic K-theory and stable smooth pseudoisotopy of discs,” invited talk at Sheffield Homotopy Miniconference, Sheffield University, May 1997.

“Two-primary algebraic K-theory of spaces and related spaces of symmetries of manifolds,” main talk at A.M.S. Summer Research Conference in Algebraic K-Theory, Seattle, July 1997.

“Two-primary algebraic K-theory of rings of integers in number fields,” main talk at Symposium on Algebraic K-Theory and Applications, ICTP, Trieste, September 1997.

“Two-primary algebraic K-theory of pointed spaces,” main talk at British Topology Meeting, Swansea, April 1999.

“Algebraic K-theory of topological K-theory,” invited talk at the conference ‘Topology, geometry and algebra: interactions and new directions’, at Stanford University, August 1999.

Scientific Talks

“On Torus Knots,” talk in seminar on manifolds, University of Oslo, 1984.

“Equivariant stable homotopy theory and the Segal conjecture,” talk in Topology seminar, University of Oslo, 1989.

“Rank Filtration in Algebraic K-theory,” talks at Algebraic topology seminar, MSRI, Berkeley, 1990, at K-theory seminar, Universität Bielefeld, 1991, and at Topology seminar, Aarhus University, 1991.

“Functor Calculus,” series of 4 lectures at the University of Oslo, 1990.

“Equivariant K-theory of Spaces,” talk in Topology seminar, Aarhus University, 1991.

“Bökstedt and Madsen’s computations of T C(Z, p),” talk in K-theory seminar, University of Oslo, June 1992.

“Filtering the spectrum K(Z) by rank,” talk in Topology meeting, Oberwolfach, September 1992.

“Approximating K_{∗}(Z) through degree five,” talk in K-theory seminar, Université Paris 7, October 1992.

“K_{4}(Z) is the trivial group,” talks at Topology seminar, University of Oslo, February 1993, at Topology seminar, Aarhus University, May 1993, at K-theory seminar, Université Paris 7, May 1993, at K-theory seminar, Universität Bielefeld, June 1993, at Algebraic K-theory meeting, Oberwolfach, June 1993, and at Colloquium, Universität Wuppertal, September 1993.

“The multiplication on T HH(Z),” talk at Mittag–Leffler institute, Stockholm, February 1994.

“Algebraic K-theory, some applications and computations,” colloquium talk at University of Trondheim, AVH, March 1994.

“T C and K-theory of the two–adic integers,” talk at Mittag–Leffler institute, Stockholm, May 1994.

“The Hatcher–Waldhausen map is an infinite loop map,” talk at Topology seminar, University of Oslo, August 1994, and at Topology meeting, Oberwolfach, September 1994.

“Vassiliev’s knot invariants,” 3 talks in Topology seminar at University of Oslo, October 1994.

“Algebraic K-theory of the two-adic integers,” talk in Topology seminar, University of Oslo, April 1995.

“Fixed points of topological Hochschild homology at two,” talk in Topology seminar, University of Oslo, April 1995.

“The fiber of the linearization map from A(∗) to K(Z),” talks at Topology meeting, Oberwolfach, September 1995, and at the University of Chicago, April 1996.

“Algebraic K-theory of the two-adic integers,” talks at Algebraic K-theory and Homotopy Theory meeting, Oberwolfach, November 1995, at Stanford University, January 1996, at Northwestern University, April 1996, at the University of Chicago, May 1996, at M.I.T., May 1996, and at Algebraic K-theory meeting, Oberwolfach, June 1996.

“Introduction to topological cyclic homology,” two talks at the University of Chicago, April 1996.

“Two-primary algebraic K-theory of rings of integers in number fields,” talks at Topology seminar, University of Oslo, November 1996, at Topology meeting, Oberwolfach, September 1997, at Topology seminar, Aarhus University, December 1997, and at Algebraic Geometry and Number Theory seminar, Université de Paris Sud XI, Orsay, February 1998.

“The chain rule in functor calculus,” talk in Topology seminar, University of Oslo, February 1997.

“Equivariant stable homotopy theory,” series of 4 talks in Topology seminar, University of Oslo, April 1997.

“Homology and cohomology operations,” series of 3 talks in Topology seminar, University of Oslo, November 1997.

“Cohomology of the smooth Whitehead spectrum,” talks at Topology seminar, Universität Bielefeld, December 1997, at Topology seminar, University of Oslo, January 1998, at Symposium on ring spectra, K-theory and trace invariants, Oslo, March 1998 and at Barcelona conference on Algebraic Topology, Bellaterra, June 1998. “K_{4}(Z) is the trivial group, revisited,” talk at Topology seminar, University of Oslo, March 1998. “Two-primary algebraic K-theory of pointed spaces,” talks at Polynomial Functors conference, Schloß Ringberg, January 1999, at Workshop on Stable homotopy theory and algebraic K-theory, Universität Bielefeld, February 1999, and at West Coast Topology meeting, Stanford University, April 1999.

Popularized Talks

“Knuter og Uknuter,” talk for high school students, at University of Oslo, March 1993. “Vassilievs Knuteinvarianter,” talk for visiting students from Aarhus University, at University of Oslo, October 1994. “Fermats Siste Teorem,” talk for high school teachers at Allmennpedagogisk dag, University of Oslo, January 1995, and as popularized talk during Bjørnegildet, March 1995. “Algebraisk K-teori og symmetrier av mangfoldigheter,” 2 talks at Ski og Matematikk winter conference, Rondablikk, January 1998. “Knuteteori,” talk for finalists of the N.H. Abel competition, March 1998.

Contribution in Media

Radio interview in “Dagsnytt atten,” NRK P1, concerning Simon Singh’s book Fermat’s Last Theorem. October 1998.

Other Meetings

Topology seminars at the Institute of Advanced Study, Princeton, New Jersey, 1986–1989. Symposium in Automatic Groups, Princeton, 1989. Year in Algebraic Topology at MSRI, Berkeley, 1989–1990, including workshops on Homotopy Theory, Applications of Algebraic Topology to Geometry and Analysis, and K-theory. West Coast Topology Symposium, Stanford University, March 1990 and March 1996. Elliptic Generae meeting, Oberwolfach, October 1991. International K-theory Conference, Strasbourg, July 1992. First European Congress of Mathematics, Paris, July 1992. Topology meeting, Oberwolfach, September 1993 and September 1996.

Namboodiri lectures, University of Chicago, May 1996. Midwest topology seminar, University of Chicago, May 1996.

Refereeing, Evaluations

For Journal of Pure and Applied Algebra, 1994. For Mathematica Scandinavica (two papers) and for Topology, 1996. For Stichting Mathematisch Centrum, 1996. For Journal of Pure and Applied Algebra, for Topology and for Proceedings of the American Mathematical Society, 1997. For ICTP Symposium proceedings and K-Theory, 1998. Opponent at Morten Brun’s public Ph.D. defense, at Aarhus University, February 1998. For Acta Mathematica, the National Science Foundation (twice) and BCAT 1998 Conference proceedings, 1999.

Grants, Fellowships, Prizes

Kong Haakon den 7des utdannelsesstipend for norsk ungdom, 1985. Several student scholarship grants, University of Oslo, 1984–1986. Scientific stipends, University of Oslo, 1985–1993, from H. Ths. Heftyes legat, Hans E. Hurums minnefond, Professor S. A. Sexes legat (3 times), Stud. med. Morten Dedekam Harboes legat (2 times), Gustav Adolf Bruuns legat, and Professor Fredrik Petersens fond. Fulbright–Hays Fellowship for graduate studies in Mathematics in U.S.A., 1986–1987. NTNF (Royal Norwegian Council for Scientific and Industrial Research), Research fellowship for advanced studies in U.S.A., 1986–1987. Several honor fellowships from Princeton University, 1986–1990. Norway–America Association, Cornelia Glomme–Drolsom Award for Ph.D. studies in Mathematics at Princeton University, 1987–1988. NAVF (Norwegian Research Council for Science and the Humanities), Dr. stipend (Doctoral fellowship) for Ph.D. studies in Mathematics at Princeton University, 1987–1990. Professor Ingerid Dal og søster Ulrikke Greve Dals pris for humanistisk forskning, 1999.

Membership, Scientific Societies

Norwegian Mathematical Society, 1980–1981 and 1992–present. American Mathematical Society, 1986–1991.

Administrative Experience

Medlem i Matematisk Fagutvalg, University of Oslo, 1984–1985. Blæstsjef i Realistforeningen, University of Oslo, 1985–1986. Viseformann i Norsk Matematisk Forening, 1995–1999. Representant i instituttstyret, Matematisk institutt, 1997–1999. Prosjektleder for NFR-prosjekt “Algebraisk K-teori og topologi,” 1998–2000. Organized “Symposium on ring spectra, K-theory and trace invariants” at University of Oslo, March 26–28th 1998. Organized “Symposium on Algebraic K-theory and Topology” at University of Oslo, May 25–26th 1999.

Competitions

Niels Henrik Abel konkurransen. Three 1st and one 2nd price, 1981–1984. Chemistry Olympics, Timis.oara, Rumania, 1983. 3rd price. Mathematics Olympics, Prague, Czechoslovakia, 1984. 3rd price. Young Scientists competition in Norway, and International Final in Eindhoven, Holland, 1984. 1st price in both.

Publications and Preprints

J. Rognes: “On torus knots and the composition of mathematical knots,” International Baccalaureate Extended Essay, 1. prize in the Young Scientists competition, 1984. J. Rognes: “The rank filtration in algebraic K-theory,” Ph.D. thesis, Princeton University, 1990. J. Rognes: “A spectrum level rank filtration in algebraic K-theory,” Topology 31, 813–845 (1992). J. Rognes: “The Tits buildings for Z/p^{n},” preprint, 1991. J. Rognes: “Equivariant K-theory of spaces,” preprint, 1991. J. Rognes: “Approximating K_{∗}(Z) through degree five,” K-Theory 7, 175–200 (1993). J. Rognes: “Equivariant functor calculus,” preprint, 1991. J. Rognes: “Characterizing connected K-theory by homotopy groups,” Math. Proc. Camb. Phil. Soc. 114, 99–102 (1993). J. Rognes: “The Hatcher–Waldhausen map is a spectrum map,” Math. Ann. 299, 529–549 (1994).

J. Rognes: “The Segal conjecture for equivariant A-theory fails,” preprint, 1992. J. Rognes: “K_{4}(Z) is the trivial group,” preprint, 1993. J. Rognes: “The spectrum homology of JK(Z),” preprint, 1993. B. Jahren and J. Rognes: “An involutive cyclotomic trace map,” in preparation. J. R. Klein and J. Rognes: “The fiber of the linearization map A(∗) → K(Z),” Topology, 36, 829–848 (1997). J. R. Klein and J. Rognes: “A chain rule in the calculus of homotopy functors,” submitted to Geometry and Topology. J. Rognes, after M. Bökstedt: “Trace maps from the algebraic K-theory of the integers,” J. Pure Appl. Algebra, 125, 277–286 (1998). J. Rognes: “The product on topological Hochschild homology of the integers with mod four coefficients,” J. Pure Appl. Algebra, 134, 210–217 (1999). J. Rognes: “Topological cyclic homology of the integers at two,” J. Pure Appl. Algebra, 134, 219–286 (1999). J. Rognes: “Algebraic K-theory of the two-adic integers,” J. Pure Appl. Algebra, 134, 287–326 (1999). J. Rognes: “K-theory of the integers and stable pseudoisotopy of a point,” preprint, 1996. J. Rognes and C. Weibel: “Two–primary algebraic K-theory of rings of integers in number fields,” Journal of the A.M.S., 13, 1–54 (2000). J. Rognes: “Two-primary algebraic K-theory of spaces and related spaces of symmetries of manifolds,” Algebraic K-Theory, eds. W. Raskind and C. Weibel, Proc. Symp. Pure Math., vol. 67, A.M.S., 213–229 (1999). J. Rognes: “K_{4}(Z) is the trivial group,” Topology, 39, 267–281 (2000). J. Rognes and P. A. Østvær: “Two-primary algebraic K-theory of two-regular number fields,” Math. Z, 233, 251–263 (2000). J. Rognes and C. Weibel: “Étale descent for two-primary algebraic K-theory of totally imaginary number fields,” K-Theory, 16, 101–104 (1999). J. Rognes and C. Weibel: “Two-primary algebraic K-theory of curves over finite fields,” preprint, 1998. J. Rognes: “Two-primary algebraic K-theory of pointed spaces,” submitted to Acta Math., 1999. J. Rognes: “On the image of Milnor’s K_{4} in Quillen’s K_{4} for number fields,” in preparation, 1999. C. Ausoni and J. Rognes: “Algebraic K-theory of topological K-theory,” in preparation, 1999.